OpenStudy (anonymous):

Let f(x) = 1/sqrt(x+1) , find (f(x+h)-f(x))/h . Simplify as much as possible.

OpenStudy (anonymous):

lots of ugly algebra so lets go slow and take it step by step first off, what is \(f(x+h)\) in this case ?

OpenStudy (anonymous):

Would it be \[(1/\sqrt{x+1})+h ?\]

OpenStudy (anonymous):

no

OpenStudy (anonymous):

that is not \(f(x+h)\) that is \(f(x)+h\)

OpenStudy (anonymous):

where you see an \(x\) in \(f(x)=\frac{1}{\sqrt{x+1}}\) replace it by \(x+h\)

OpenStudy (anonymous):

Sorry, I'm kind of confused! Algebra is not my forte. Would the equation then be \[1/ \sqrt{x+h+1} ?\]

OpenStudy (anonymous):

yes good

OpenStudy (anonymous):

now if you don't like algebra, you are really not going to like the next part

OpenStudy (anonymous):

we have \[f(x+h)=\frac{1}{\sqrt{x+h+1}}\] so \[f(x+h)-f(x)=\frac{1}{\sqrt{x+h+1}}-\frac{1}{\sqrt{x+1}}\] and now you have to actually do the subtraction any ideas?

OpenStudy (anonymous):

Convert the two fractions to common denominators?

OpenStudy (anonymous):

by rationalizing the denominators

OpenStudy (anonymous):

yeah, but don't do too much work leave the denominator in factored form \[\frac{1}{\sqrt{x+h+1}}-\frac{1}{\sqrt{x+1}}=\frac{\sqrt{x+1}-\sqrt{x+h+1}}{\sqrt{x+h+1}\sqrt{x+1}}\]

OpenStudy (anonymous):

that wasn't too bad, that is just like \(\frac{1}{a}-\frac{1}{b}=\frac{b-a}{ab}\)

OpenStudy (anonymous):

now comes the real kicker rationalize the NUMERATOR don't multiply anything out in the denominator at all, leave it in factored form but multiply top and bottom by \(\sqrt{x+h+1}+\sqrt{x+1}\)

OpenStudy (anonymous):

ok that was a mistake, the conjugate of the numerator is \[\sqrt{x+1}+\sqrt{x+h+1}\]

OpenStudy (anonymous):

still with me or did i lose you yet?

OpenStudy (anonymous):

Kiiiiinda but I'm kind of confused as of how you'd subtract the \[\sqrt{x+h+1}\]

OpenStudy (anonymous):

messed that up \[\frac{a}{b}-\frac{c}{d}=\frac{ad-bc}{bd}\]

OpenStudy (anonymous):

so i didn't really "subtract" i just wrote it

OpenStudy (anonymous):

\[ \it's \sqrt{x+1}-\sqrt{x+h+1}/\sqrt{x+1}\sqrt{x+h+1}\]

OpenStudy (anonymous):

\[\frac{1}{\sqrt{x+h+1}}-\frac{1}{\sqrt{x+1}}=\frac{\sqrt{x+1}-\sqrt{x+h+1}}{\sqrt{x+h+1}\sqrt{x+1}}\] is just subtraction notice we did not multiply anything out or combine like terms or any algebra at all,

OpenStudy (anonymous):

Is that the final answer? i don't thinkyou can simplify that any more

OpenStudy (anonymous):

oh not even close

OpenStudy (anonymous):

\[\frac{\sqrt{x+1}-\sqrt{x+h+1}}{\sqrt{x+h+1}\sqrt{x+1}}\times \frac{\sqrt{x+1}+\sqrt{x+h+1}}{\sqrt{x+1}+\sqrt{x+h+1}}\]\]

OpenStudy (anonymous):

leave the denominator in factored form, but a miracle occurs in the numerator

OpenStudy (anonymous):

this is the rationalizing the numerator part

OpenStudy (anonymous):

\(a-b)(a+b)=a^2-b^2\) so the numerator gives \[x+1-(x+h+1)=-h\]

OpenStudy (anonymous):

the \[\sqrt{x+1}\] cancels out right?

OpenStudy (anonymous):

Oh, I see!

OpenStudy (anonymous):

yeah both square roots go you have to only be careful with the subtraction, make sure you realize the numerator is \(-h\) and not \(h\) only one more step...

OpenStudy (anonymous):

Yay! Getting close :) So at this point, the equation is \[ \frac{{-h} }{\sqrt{x+1}+\sqrt{x+h+1} } \] isn't it?

OpenStudy (anonymous):

you are missing a huge part of the denominator

OpenStudy (anonymous):

\[\frac{\sqrt{x+1}-\sqrt{x+h+1}}{\sqrt{x+h+1}\sqrt{x+1}}\times \frac{\sqrt{x+1}+\sqrt{x+h+1}}{\sqrt{x+1}+\sqrt{x+h+1}}\] \[=\frac{-h}{(\sqrt{x+h+1}\sqrt{x+1})(\sqrt{x+1}+\sqrt{x+h+1}}\]

OpenStudy (anonymous):

don't even think about multiplying out in the denominator, it will be a huge mess just leave it

OpenStudy (anonymous):

That's the final answer?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

remember that \(h\) in the denominator we have been ignoring this whole time? \[\frac{f(x+h)-f(x)}{\color{red}h}\]

OpenStudy (anonymous):

it cancels with the \(-h\) in the numerator, leaving a \(-1\) up top

OpenStudy (joftheworld):

@satellite73 can you help alittle more when you're done ?

OpenStudy (anonymous):

\[\frac{-1}{(\sqrt{x+h+1}\sqrt{x+1})(\sqrt{x+1}+\sqrt{x+h+1}}\] is the final answer

OpenStudy (anonymous):

Thanks! But, wouldn't the numerator be -1-\[\frac{ 1 }{ \sqrt{x+1} }\] ? because of the extra -f(x) after the f(x+h)? Or did we already do that and I just forgot?

OpenStudy (anonymous):

numerator is \(-1\) for sure, because before cancelling it was \(-h\) the step to making it \(-h\) was \[(\sqrt{x+1}-\sqrt{x+h+1})(\sqrt{x+1}+\sqrt{x+h+1})=x+1-(x+h+1)=-h\]

OpenStudy (anonymous):

Oh, alright! I see now. Thanks for you help and patience. :) Sorry, that was for a summer packet for precal and they wanted us to figure how to do these problems by ourselves. Thanks again!

OpenStudy (anonymous):

yw this was not that enlightening, just a lot of algebra in calc you will repeat this process, then replace \(h\) by \(0\) to get the "derivative" fortunately there are lots of shortcut methods so you wont have to do this much